Mar 26, 2013 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  1. Monitoring
    • Supply a draft specification for preparing the surfaces of the FCAL plexiglass sheet (Athens)
  2. SiPM
    • After summary of SiPM parameters, provide data to Hamamatsu for comments (USM)

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Monitoring
  4. FCAL
  5. BCAL
    1. USM Report File:20130326.pdf
    2. SiPM Measurements at Regina
  6. Reconstruction and Simulation
  7. TDR update
  8. Calibration of Calorimeters
  9. Any other business


Attendees:Elton, Tim, Yi, David, Beni (JLab); Kei, Matt (IU); Zisis, Andrei (UofR); Will, Hayk, Alam, Elias (USM); Christina, Stratos, Pavlos (Athens)

  1. Announcements
    • News media event in Valparaiso will happen this morning.
  2. Action Items
  3. Monitoring
    • Almost all material in hand
    • LEDs have not yet arrived. (Note: Zisis said that they were shipped 8 days ago and expect arrival in a few days)
    • Boards will start soon and be tested for functionality. Athens will check pulse sizes relative to requirement.
    • Stratos will work on the plexiglass specs. Once testing is completed, a piece of the prototype in Athens will be shipped to JLab as a sample.
  4. FCAL
    • Stacking is proceeding very well. Expect completion at the end of the week or early next week,
    • Zisis: Plan a photo shoot for the final block. (Elton will ask Greg Adams to take photo).
    • Bases are being completed slowly.
    • Drawing package on the dark box has been completed. Welds on the aluminum structure, which must be derated, have been replaced by bolted connections.
    • Tim: Has seen the model. Drawing numbers were issued yesterday. Drawings will be entered into the JLab system before installation.
    • Kei: Draft on NIM paper from beam test has been circulating. Matt, Kei and Elton stayed after the meeting to go over recent comments.
  5. BCAL
    1. USM Report File:20130326.pdf (Note: As the report was started, USM was disconnected and was unable to reconnect for the rest of the meeting).
      • All measurements have been completed at USM. The next-to-last box (640 sensors) should have arrived yesterday at JLab. The last shipment (370 sensors) was sent out yesterday.
      • The 3 non-conforming sensors will be shipped to JLab as part of the last shipment of 370 sensors. Jlab will return them to Hamamatsu for replacement.
      • Final testing on the forty original sensors on loan to USM (from the first article) will be completed in the next week and they should be shipped back in about a week. They will make linearity measurements on these sensors.
    2. SiPM Measurements at Regina
      • Good progress has been made on PDE measurements
      • Recent measurements are consistent with original data from Mernoosh. SiPM and diode are being flipped in position to check consistency.
      • Plan to complete 7 or 8 sensors by the end of this week, the remaining will be measured next week. Measurements are taken at 0.9 V above breakdown. If time permits, other measurements will be taken at different biases.
    3. JLab production (Elton)
      • Production is proceeding at a steady pace. So far 26 modules have been completed and tested
      • Modules have been moved to the hall and the staff has started mounting them on the cradle. Five modules have been mounted. A couple of modules were found with a broken light guide on one of the top corners. We do not know that have been bumped, but they are the most vulnerable. They have been reglued in place.
      • Module.... was found to have some reflecting stripes in some of the light guides in the middle of the module face. The concern was that the light guides may be coming unglued. However, after discussion and inspection by several people, we believe that the stripes are due to reflections of glue squeezed up between light guides.
      • We will try to develop a portable testing setup that can help evaluate quality of optical contacts. Zisis: Will send Elton a description of the test setup that was used early on to check machined modules. Andrei: May be difficult to be sensitive to small changes in transmission.
      • Shaun: Developing a program to check the cosmic-ray runs. Draft version sent to Zisis and Andrei to request feedback.
  6. Reconstruction and Simulation
    • Did not meet last week
    • Will is looking at existing code, and considering how it needs to be updated to match the latest simulation updates from David.
    • Andrei has sent Will a description of a methodology for iterative time-walk corrections. Also checking literature for algorithms that can separate two merging clusters.
  7. TDR update
    • Zisis: With the fiber NIM paper approved, time will open up to get back to the TDR early next month. Expects to have a draft for feedback relatively soon.
  8. Calibration of Calorimeters
  9. Any other business